# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ 4.2 Risk-coping Mechanisms
Preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, become smaller when I focus on the semi-monthly frequency. Specifically, food expenditure is likely to be insured within at least half a month. The workers might have dealt with their shorter-run shocks relatively well but could not cope with the longer-run shocks. In this section, I test the mechanisms behind their smoothing behavior using several variables on the risk-coping devices.
@@
To test the roles of risk-coping strategies, I employ the empirical specification suggested by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is characterized as follows:
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I prefer to use the head’s income instead of the disposable income in equation 2 to avoid the potential endogeneity. [^64] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results are largely unchanged if I use the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I prefer to use the conservative setting to deal with potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. As for $r_{i,t}$, I consider several variables to overview various risk-coping strategies. These include the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable is defined as the difference between income and expenses. The estimates of $\kappa$ are therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also consider the earnings of all the family members except for the head and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales may play a role in mitigating the idiosyncratic shocks as alternative risk-coping strategies (Udry 1995; Blundell et al. 2016). [^65] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focuses on the households living in Tsukishima throughout the sample period, I can discard addressing the ex-ante risk-coping strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
@@ Results
This table shows the results of equation 2: the family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
@@
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
Table 5 presents the results. Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively. Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of assets.
@@
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen. The estimate from the monthly series is similar ($0.042$), albeit weakly statistically significant (Column (2)). This is basically consistent with the historical fact that many workers had savings accounts both in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitude is small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
@@
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source in the category is the mutual loan associations (mujin). As explained in Section 2.2, these associations’ rules were unsuitable for compensating the idiosyncratic shocks. A suggestive case is that a THBS household won a lottery and got $100$ yen in a monthly cell, which is clearly independent of the worker’s earnings. [^66] The OISR, Archives of the Tsukishima Survey (THBS #5, 17th May 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the incomes from mujin measured in the budget books are less than one yen, which is not the case on average.
@@
For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns. Section 2.2 shows that pawnshop was the popular lending institution for low-income households. Although THBS households occasionally used pawnshops, they never used them frequently. [^67] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This makes sense because skilled factory workers can not classified as the low-income working class.
@@
The estimated coefficient for the net income from credit purchases is negative and statistically significant in Column (1). The estimate suggests that a one yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen. This magnitude is larger among the other categories. According to the THBS budget books, the credit purchases were mostly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). Since rice, a staple food, is usually an expensive item, credit purchases might have played an important role among these households. The credit purchases were occasionally used in the call doctor and pharmacy. This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4). While this provides suggestive evidence that credit purchase is the primary risk-coping strategy, Column (2) shows that the credit was less likely to work in the longer run. This seems to be reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5). Instead, the result for the net savings category suggests that precautionary savings have helped mitigate income declines lasting at least one month due to idiosyncratic shocks. Given the magnitudes, however, it is important to note that it could only partially compensate for the idiosyncratic shocks.
@@
The estimates for the net income from gifts are negative and statistically significant in Column (1). The budget books tell that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is calculated to be roughly $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns are close to zero and statistically insignificant. Miyoshi (1989, pp. 383–386) documented that the wives regularly engaged in side jobs rather than worked temporarily to compensate for the income loss. In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^68] The OISR, Archives of the Tsukishima Survey (THBS#1, 13th January 1919). In the final row in Panel B, I find that selling miscellaneous assets such as newspapers, bottles, and clocks did not work as a risk-coping strategy.
@@
To summarize, I find that credit purchases have played a role in mitigating the idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these were basically available for short-term, say at most half a month. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions, which is indeed consistent with the findings on the partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding, which is not explicitly described in the historical sources, is the possibility that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
4.3 Event-Study Analysis: Response to the Advers Health Shocks
@@
The foregoing analyses used the within variations in the earnings for estimating the income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the shocks, it does not offer a clear view of the pathways behind the risk-coping behaviors against the adverse idiosyncratic shocks. To clarify this point, I next focus on the temporary illness of the heads that cause temporary income losses.
@@
I begin my analysis by testing whether the illness of the head impacted the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly death counts from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books suggest that the family members had caught the flu during this period. Five households experienced the illness during the sample period, and the illness affecting the heads was concentrated between February and May 1919, when flu exposure was most prevalent. [^69] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. In addition, the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$), which is consistent with the fact that Type A influenza causes fever for less than $6$ days (Uyeki et al. 2022). Therefore, it is plausible to see that the influenza epidemics largely led to illnesses during this period, providing a quasi-experimental setting in analyzing to what extent the adverse idiosyncratic shock driven by illness influences the households’ consumption behaviors.
@@
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings shall mainly be influenced by the number of days suffering illness in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on his earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in the illness decreases the head’s semi-monthly income by roughly $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). Thus, this reflects the reductions in their salary due to the absence: Miyoshi (1989, pp. 383; 388) highlighted that income was accidentally reduced due to absence when the head was infected with a disease. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^70] Given that the average number of days suffering the illness is roughly $2.0$, the temporary health shock loses more than $5$ yen in that month, accounting for roughly $10$% of the average monthly income. Although this is a rough estimate and the losses shall depend on the severity of the shocks, my result at least shows that the household income had reacted to the adverse health shock among the factory workers’ households.
@@
Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including eather of the two variables does not change the finding on the neglisible impacts of the health shocks on the other family members.
Following the evidence, I characterize equation 1 as a structural consumption smoothing equation by using the first-stage reduced form equation as follows:
@@
where $h_{i,t-1}$ is the number of days of head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. [^71] As noted in Section 2.3, the factories did not offer compensation for their sickness leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916). Even in the largest factory in Tsukishima, the Ishikawajim Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930). In fact, Miyoshi (1989) documented the issue that the head’s disease directly causes income loss due to the absence. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition basically hold under this just-identified system.
@@
Smoothing Consumption against the Adverse Health Shocks
@@
Column 1 shows the results of equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household fixed effects, and time fixed effects. The estimated coefficients on the health shock variable are listed in the sub-column named “Coef.”.
@@
Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficients on the disposable income are listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 in the spirit of Cochrane (1991). Column (1) in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels. Next, I explicitly model the structure behind the relationship between the health shocks and earnings using the reduced-form equation of 3. If the negative impacts of the health shocks on consumption were derived from the income losses, the result from the structural estimation approach is similar to that in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. As shown, the results are materially similar to the reduced-form results in Column (1). Importantly, the estimates are systematically larger than those reported in Table 4. The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$. While this result partly reflects the bias reduction under the instrumental variable estimation, it is also important to remember that this system is designed to reduce the bias by specifying the channel from the adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to the rare health shock events. [^72] In other words, the furniture may have been purchased only when the households had extra income. This may partly reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^73] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on the furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applicable to the transportation category because family travel is not a frequent event and thus, unlikely to coincide with the timing of illness. In contrast, the entertainment expenses are still likely to be cut in case of illness, which may reflect the curtailment of tobacco, theater-going, and the allowance for heads. This may be another explanation for the difference in the results from my baseline results in Table 4.
While the estimates for clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on the rare health shock events, the findings are materially similar to my baseline results: both categories are sensitive to the shocks as those are dispensable categories.
Results: How did the households compensate for the losses from the ill?

### v10

@@ 4.2 Risk-coping Mechanisms
The preceding results show that the skilled workers’ households in Tsukishima were less likely to smooth their consumption. The estimated elasticities, however, became smaller when I focused on the semi-monthly frequency. Specifically, food expenditure was likely to be insured within at least half a month. The workers might have coped with their shorter-run shocks relatively well but could not manage longer-run shocks. In this section, I test the mechanisms behind their smoothing behavior using several variables in the risk-coping strategies.
@@
To test the roles of risk-coping strategies, I employed the empirical specification suggested by Fafchamps and Lund (2003). For household $i$ at time $t$, the specification is characterized as follows:
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ is the household fixed effect, $\zeta_{t}$ is the time fixed effect, and $\epsilon_{i,t}$ is a random error term. Online Appendix C.1 summarizes the derivation of this specification. I used the head’s income instead of the disposable income in Equation 2 to avoid the potential endogeneity. [^60] For instance, the household’s total income is the outcome of the intra-household responses to the shocks, such as the labor-supply adjustments. I confirmed that the results were largely unchanged if I used the disposable income as the idiosyncratic income shock variable instead of the head’s income (not reported). This is consistent with the fact that the labor-supply adjustments did not occur among the THBS households (Table 5). Despite this, I preferred to use the conservative setting to manage potential biases and to set a comparable setting to the structural equation used in Section 4.3. See Asdrubali et al. (2020) for a recent study using the head’s income as the shock variable in testing the full risk-sharing. Regarding $r_{i,t}$, I considered several variables to overview various risk-coping strategies. These included the net income on savings, insurance, borrowing, credit purchases, and gifts. Each net variable was defined as the difference between income and expenses. The estimates of $\kappa$ were therefore expected to be negative since net income becomes negative in good times but positive in unfavorable times. I also considered the earnings of the family members (except for the head) and the sales of miscellaneous goods because both intra-household labor supply adjustments and asset sales might alleviate idiosyncratic shocks as alternative risk-mitigation strategies (Udry 1995; Blundell et al. 2016). [^61] Rosenzweig and Stark (1989) show that marriage cum migration in South Indian villages contributed to reducing the variability of household food consumption. Since the THBS focused on the households living in Tsukishima throughout the sample period, I did not need to address the ex-ante risk-mitigation strategy through migration. The estimated coefficients for the sales and labor supply adjustment variables may be negative, reflecting additional earnings when the head’s income is reduced.
@@ Results
This table shows the results of equation 2. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
@@
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@
Table 5 presents the results. Columns (1) and (2) summarize the results for the semi-monthly and adjusted monthly panel datasets, respectively. Panel A shows the results for the net income variables, whereas Panel B shows those for the labor supply adjustments and sales of small assets.
@@
The estimated coefficient for the net savings is negative and statistically significant in Column (1). The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from savings by $0.043$ yen. The estimate from the adjusted monthly series is similar ($0.035$), albeit statistically insignificant (Column (2)). This is consistent with the historical fact that many workers had savings accounts in postal savings and savings banks at that time (Section 2.3). However, the estimated magnitude was small, suggesting that the withdrawals from savings did not compensate for the earnings losses.
@@
The estimates for the insurance category are close to zero and statistically insignificant in both columns. The budget books indicate that the primary source in this category was the mutual loan associations (mujin). As explained in Section 2.2, these associations’ rules were unsuitable for compensating for the idiosyncratic shocks. For example, a THBS household won a lottery and received $100$ yen in a monthly cell, which is independent of the worker’s earnings. [^62] The OISR, Archives of the Tsukishima Survey (THBS #5, May 17, 1919). The bidding rule may be used to support a specific member in the association if members have a common belief in the aid. However, most of the income from mujin measured in the budget books was less than one yen, which is not the case on average.
@@
For the net borrowing category, the estimated coefficients are close to zero and statistically insignificant in both columns. Section 2.2 shows that pawnshops were a popular lending institution for low-income households. Although THBS households occasionally used pawnshops, this was not frequent. [^63] The OISR, Archives of the Tsukishima Survey (THBS #2; 19; 22). This is understandable because skilled factory workers were not classified as low-income working class at that time.
@@
The estimated coefficient for net income from credit purchases is negative and strongly statistically significant in Column (1). The estimate suggests that a one-yen decrease (increase) in the head’s earnings increases (decreases) the net income from credit purchases by $0.13$ yen. This magnitude is the largest in all the categories. According to the THBS budget books, the credit purchases were mainly used to buy rice, fish, vegetables, and fuels (firewood and charcoal). As rice—a staple food—is usually an expensive item, credit purchases might have played an important role among these households. In addition, credit purchase was occasionally used for the call doctor and pharmacy. This is consistent with my baseline result for medical expenses in the semi-monthly series (Column (1) in Table 4). While these results suggest that credit purchase was the primary risk-mitigation strategy, Column (2) shows that credit was no longer effective in the long term. This is reasonable because borrowers could make monthly or semi-monthly payments but did not borrow for more than a month (Section 5). Given the estimated magnitude ($0.13$), however, it is noteworthy that the credit could only partially compensate for the idiosyncratic shocks.
@@
The estimates for the gifts category are negative and statistically significant in Column (1). The budget books show that the gifts included seasonal gifts, celebration money, and return gifts. Among them, the return gifts can be regarded as mutual aid from personal networks, which may be reflected in this result. According to an official report of Tokyo, however, the share of average monthly gifts in total income was approximately $3$% (Social Affairs Division 1925, pp. 58-59). In my adjusted monthly panels, this share is calculated to be approximately $4$%. In this light, while the gifts comprised a few percentage points of the monthly income, they may not have provided sufficient aid for the households facing economic hardship. In fact, the estimate is no longer statistically significant in the adjusted monthly series (Column (2)).
@@
Finally, Panel B tests the roles of the intra-household labor supply adjustment and sales of miscellaneous assets. The estimated coefficients in both columns of the first row are close to zero and statistically insignificant. Miyoshi (1989, pp. 383–386) documented that the wives regularly engaged in side jobs rather than working temporarily to compensate for the income loss. In fact, although sewing was the most popular side job among THBS wives, their earnings were too small to replace the lost income from their breadwinning husbands. [^64] The OISR, Archives of the Tsukishima Survey (THBS#1, January 13, 1919). In the second row, I show that selling miscellaneous assets such as newspapers, bottles, and clocks was not an effective risk-mitigation strategy.
@@
To summarize, the evidence suggests that credit purchases mitigated idiosyncratic shocks. The results also indicate that the savings institutions and personal networks might have been used as risk-coping institutions among the workers. However, these were available for the short term —at most half a month. The estimated magnitudes indicate that the workers could only partially compensate for the temporary losses of income by using these institutions. This supports the findings on partial risk-sharing in Section 4.1. Overall, the results are aligned with the descriptives from the historical documents on the financial institutions available in Tokyo City (Section 2.3). An interesting finding—which is not explicitly described in the historical sources—is that the credit purchases in the local society and the informal gifts from personal networks might have provided some aid to the factory workers.
4.3 Event-Study Analysis: Response to Adverse Health Shocks
@@
The foregoing analyses used the within variations in the earnings to estimate income elasticity, meaning that the identification depends both on the adverse and favorable idiosyncratic shocks. While it is useful to illustrate the overall responses of the households to the shocks, it does not reveal the pathways behind the risk-coping strategies. Next, I focus on temporary illness of household heads that could cause temporary income losses.
@@
I began my analysis by testing whether the illness of the head impacted the earnings. To do so, I decided the exact timing and duration of illnesses of the family members using the Kazoku idō hyō (Section 3.4). Figure 5 shows the number of monthly deaths from influenza (ryūkōsei kanbō) in Tokyo between 1918 and 1920. There appears to have been a moderate spike in early 1919, which is partially captured in the THBS period. In fact, a few THBS budget books indicate that the family members contracted influenza during this period. Five households experienced the illness during the sample period, and illness affecting the heads was concentrated between February and May 1919, when influenza exposure was most prevalent. [^65] Among four cells with health shocks on the heads, two are observed in February, during the peak of the epidemics (Figure 5), and the others are in March and May. In addition, the average duration for the heads is $5.8$ days (Std. Dev. $=3.6$), which is consistent with the fact that Type A influenza caused fever for less than six days (Uyeki et al. 2022). Therefore, it is plausible that influenza was responsible for illnesses during this period, providing a quasi-experimental setting for analyzing to what extent the adverse idiosyncratic shock of illness influences the households’ consumption strategies.
@@
Earnings in a semi-month are based on the work done during the previous semi-month. This means that the head’s earnings were influenced by the number of illness days in the previous semi-month cell. Table 6 presents the results from the regression of the head’s earnings on the number of illness days in the previous semi-monthly cell. While the instantaneous health shock has little impact on the head’s earnings (Column 1), the lagged health shock has a statistically significant effect on their earnings (Column 2). This result is unchanged if I include both variables in the same regression in Column 3, supporting the evidence that the health shocks in the previous cell reduced the earnings of the head. The estimate in Column (2) suggests that a one-day increase in illness decreases the head’s semi-monthly income by approximately $2.4$ yen. Notably, this is close to the average daily wage rate of skilled factory workers (Online Appendix B.6). This reflects the reductions in their salary due to absence: Miyoshi (1989, pp. 383; 388) highlighted that income was reduced due to absence when the head was infected with a disease. Finally, Column (4) adds the health shocks of the other family members into the specification of Column (3). The estimated coefficients on these placebo variables are statistically insignificant, supporting the robustness of the result in Column (3). These results confirm that the head who suffered illness lost earnings. [^66] Given that the average number of illness days was approximately $2.0$, the temporary health shock losses were more than five yen in that month, accounting for approximately $10$% of the average monthly income. Although this is a rough estimate and the losses depend on the severity of the shocks, my result shows that the household income responded to the adverse health shock of illness among the factory workers’ households.
@@
Notes: This table shows the results from the regressions of the head’s earnings on the health shock variables using the semi-monthly panel dataset. A household with two semi-month observations is excluded from the sample in all the regressions because the lagged health shock variables are included. The family size variable, household fixed effects, and time-fixed effects are included in all the regressions. “Health shocks on the head” refers to the number of days that the head suffered illness. “Health shocks on the others” refers to the number of days that family members, excluding the head, were ill. “[level]” indicates the variable of the head’s earnings in the same semi-month cell. “[lag]” indicates the lagged variable (i.e., the shocks measured in the previous semi-month cell). Although the correlation coefficient between the level and lag of the health shock variables for the other family members is substantially high (Pearson’s coefficient $=0.99$), including either of the two variables does not change the finding on the negligible impacts of the health shocks on the other family members.
Following the evidence, I characterize Equation 1 as a structural consumption-smoothing equation by using the first-stage reduced form equation as follows:
@@
where $h_{i,t-1}$ is the number of days of the head’s illness in the previous semi-month cell, $\varpi_{i}$ is the household fixed effect, $\tau_{t}$ is the semi-month fixed effect, and $e_{i,t}$ is a random error term. For the risk-sharing mechanisms analysis under equation 2, the log-transformed disposable income in equation 3 ($y_{i,t}$) is replaced with the head’s earnings ($\tilde{y}_{i,t}$). The health shock variable is plausibly exogenous because the illnesses were from the influenza epidemics, as explained above. [^67] As noted in Section 2.3, the factories did not offer compensation for sick leave because the Factory Act did not specify benefits for workers’ illness (Study Group of the Factory Act 1916). Even in the largest factory in Tsukishima, the Ishikawajima Shipyard, any descriptions of the influenza epidemics are documented in the company’s history (Arai 1930). In fact, Miyoshi (1989) documented the issue that the head’s disease directly caused income loss due to the absence at that time. In addition, there is a correlation between the health shock variable and the head’s earnings at a statistically valid level (Table 6). Both support the evidence that the exogeneity of the instrument and relevance condition hold under this just-identified system.
@@
I first consider a reduced-form specification using the head’s illness as the exposure variable in equation 1 following Cochrane (1991). Column (1) in Table 7 shows the results. Overall, the coefficients on the exposure variable are negative, supporting evidence that adverse health shocks on the heads reduce consumption levels. Next, I explicitly model the structure underlying the relationship between the health shocks and earnings using the reduced-form equation of (3). If the negative impacts of the health shocks on consumption were derived from the income losses, the results from the structural estimation approach correspond to those in Column (1). The estimates from the structural equation are listed in Column (2) of Table 7. As shown, the results are congruent with the reduced-form results in Column (1). Importantly, the estimates are systematically larger than those reported in Table 4. The estimates for total and food consumption categories are $0.543$ and $0.355$, respectively, whereas the baseline estimates are $0.358$ and $0.136$. While this result partly reflects the bias reduction under the instrumental variable estimation, it is also noteworthy that this system is designed to reduce the bias by specifying the channel from adverse (instead of favorable) income shocks. For example, expenses on the allowance for children could be easily cut during unfavorable times but may not respond to favorable shocks. Similarly, the expenses for furniture are likely to be cut, but the households might not buy those frequently, meaning that there are no sufficient cases for reacting to rare health shock events. [^68] In other words, furniture may have only been purchased when the households had extra income. This may partly be reflected in the smaller value of the first-stage $F$-statistics in the furniture subcategory. [^69] Note that the logarithmic transformation leaves the cells with expenses on the furniture. Thus, the reduced-form equation captures the relationship between the health shocks and earnings in the cells with any expenses on furniture. While the first-stage $F$-statitics is $35.7$ in the linear-linear model, the second-stage result is robust against the functional form assumption. This may reflect the asymptotic theoretical advantage in the panel data model under nearly weak instruments case (Staiger and Stock 1997; Cai et al. 2012; Baltagi et al. 2012). The same argument is applicable to the housing category. A similar interpretation can be applied to the transportation category because family travel was not a frequent event and thus, unlikely to coincide with the timing of illness. By contrast, the entertainment expenses were still likely to be cut in case of illness, which might reflect the curtailment of tobacco, theater-going, and the allowance for heads. This may be another explanation for the difference in the results from my baseline results in Table 4.
While the estimates for the clothes and gifts categories now exceed one, the null of equality to the population mean of one is not statistically rejected for both cases. Although the large estimates may be partly due to the operation under the instrument based on rare health shock events, the findings are materially like my baseline results—both categories are sensitive to the shocks as are those of the dispensable categories.
@@
Smoothing Consumption against Adverse Health Shocks
@@
Column 1 shows the results of Equation 1, in which the log-transformed disposable income is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell). The dependent variables are the 11 measures of log-transformed consumption, and all the regressions include the family size control, household-fixed effects, and time-fixed effects. The estimated coefficient on the health shock variable is listed in the sub-column named “Coef.”.
@@
Column 2 shows the results of equation 1, in which the log-transformed disposable income is regarded as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) for the first-stage reduced-form equation (3). The estimated coefficient of the disposable income is listed in the sub-column named “Coef.”. The first-stage $F$-statistics on the health shock variable are listed in the final column.
Results: How did the Households Compensate for Losses from illness?
I tested the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results in the same manner as Table 5. The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating strategies against the shocks. The estimate for the credit purchase category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for approximately half of their average daily wages ($1.29/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable. A similar but clearer result is obtained—the estimate for the credit purchase category is approximately four times greater than that reported in Column (1) of Table 5. It indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.707070This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1). The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-mitigation strategy for dealing with health shocks. Since withdrawing and receiving gifts to use the money at the retailers takes more time (steps), the households likely prefer to use credit purchases.
Overall, the event-study analysis using adverse health shocks confirms the baseline findings and provides a sharper illustration of consumption-smoothing strategies in the case of adverse idiosyncratic shock events. My results show that the unique available response to idiosyncratic income shocks among the THBS households was to use credit purchases. The available compensation via credit purchases accounts for approximately half of their temporary income losses, which partially contributes to smoothing their food consumption in the short term.